What is the focal length of a makeup mirror that produces a magnification of 2.39 when a person’s face is 11 cm away?

answer to 3 decimal places


1
Expert's answer
2021-06-03T13:51:40-0400

Given,

Magnification of the mirror "(m)=2.39"

Distance between the face and mirror "(u)=11cm"

Let the focal length of the mirror be f.

"m=\frac{-v}{u}"

"\Rightarrow v=-11\times 2.39 cm"

"=-26.29cm"

Now, applying the mirror formula,

"\frac{1}{f}=\frac{1}{v}+\frac{1}{u}"

Now, substituting the values,

"\frac{1}{f}=\frac{-1}{26.29}+\frac{1}{11}"


"\Rightarrow \frac{1}{f}=\frac{-11+26.29}{11\times 26.29}"


"\Rightarrow \frac{1}{f}=\frac{15.29}{289.19}"


"\Rightarrow f = \frac{289.19}{15.29}"


"f=18.913cm"